# Why Smishing Works

### v2

@@ Abstract.
Abstract With the booming popularity of smartphones, threats related to these devices are increasingly on the rise. Smishing, a combination of SMS (Short Message Service) and phishing has emerged as a treacherous cyber threat used by malicious actors to deceive users, aiming to steal sensitive information, money or install malware on their mobile devices. Despite the increase in smishing attacks in recent years, there are very few studies aimed at understanding the factors that contribute to a user’s ability to differentiate real from fake messages.
To address this gap in knowledge, we have conducted an online survey on smishing detection with 214 participants. In this study, we presented them with 16 SMS screenshots and evaluated how different factors affect their decision making process in smishing detection. Next, we conducted a follow-up survey to garner information on the participants’ security attitudes, behavior and knowledge. Our results highlighted that attention and security behavioral scores had a significant impact on participants’ accuracy in identifying smishing messages. Interestingly, we found that participants had more difficulty identifying real messages from fake ones, with an accuracy of 65.6% with fake messages and 44.6% with real messages. Our study is crucial in developing proactive strategies to encounter and mitigate smishing attacks. By understanding what factors influence smishing detection, we aim to bolster users’ resilience against such threats and create a safer digital environment for all.
Smishing, Phishing, Usable Security
@@ 1. Introduction
Phishing stands out as one of the most prevalent social engineering attacks in cybersecurity, with Cisco reporting that phishing accounts for 90% of data breaches in the U.S. (cisco). These attacks collectively lead to billions of dollars in losses annually (whitehouse). Attackers are now extending their efforts into the mobile domain, introducing ransomware, spyware, or adware onto victims’ mobile devices through phishing (bitdefender). This malicious software can provide attackers access to sensitive information, including passwords, credit card details, location data, and social security numbers, resulting in significant financial damage (nagunwa2014behind).
@@
A survey by Openmarket reveals that 75% of Millennials prefer texting over phone calls, and 83% read SMS messages within 90 seconds of receipt (eztextingstats). With the ubiquity of smartphones in daily life, attackers are targeting billions of users (statista). RoboKiller’s data indicates a surge in spam texts, with over 225 billion sent in 2022, leading to losses exceeding 22 billion dollars (robocall). This represents a 157% increase from the previous year and a staggering 307% rise since 2020. According to the Federal Trade Commission, in 2022, consumers reported losing \$330 million to text message scams, which is more than double the amount reported in 2021 (FTC300million).
@@
Numerous classical studies (dhamija2006phishing; sheng2010falls; alsharnouby2015phishing) have sought to understand what makes users vulnerable to email-based phishing attacks. However, research in the realm of smishing remains limited. In a recent study by Rahman et al., (rahman2023users), an empirical examination of smishing was carried out. They sent eight fake SMSes to 265 users to gauge the effectiveness of smishing attacks. Their findings revealed that up to 16.92% of the recipients could have fallen victim to these attacks. In a separate study on smishing, Blancaflor et al. (blancaflor2021let) discovered that one in 24 targeted users clicked on the phishing URLs within the SMS. These investigations delved into the number of fake SMSes that reached participants, considering that smishing messages could be intercepted by SMS gateways, mobile carriers, or anti-smishing apps on users’ devices (timko2023commercial).
@@
To gain a comprehensive understanding of the issue, a more controlled experiment is needed—one that evaluates users’ ability to detect smishing and examines their interactions with both real and fake messages. While previous studies (rahman2023users; blancaflor2021let) have investigated smishing from the perspective of whether a user actively becomes a victim, we intend to approach the topic from the users’ perception of the messages. We explore what triggers users’ suspicion that a message may be fraudulent, thus causing them to lose their comfort or willingness to interact with it. Additionally, there is nuance between a user actively interacting with a phishing message and whether they believe it is legitimate. End users may perceive a phishing message as legitimate or a legitimate message as fraudulent, while still refraining from interacting with it. By investigating this nuance, we can build a broader picture of smishing and, in turn, help focus security experts and developers on the factors causing end users to make these classification errors.
@@
In this paper, our objective is to bridge the existing gaps by undertaking a comprehensive online study. This will offer a more profound insight into the influence of messaging content and user characteristics on interaction rates, comfort levels, and judgement between real and fake messages. In the main part of the study, we carried out an online user survey focused on smishing detection. To draw correlations with various factors, we incorporated a follow-up post-survey that encompasses several standardized metrics.
@@
    We designed and conducted an online survey with 214 participants to gain insights into how different user and message factors affect users’ likelihood to interact, comfort levels with interacting and their judgement of real and fake messages.
@@ 2. Related Works
Research by Dhamija et al. (dhamija2006phishing) marked the first investigation into understanding the most effective strategies for deceiving victims in phishing attacks. The study revealed that people consistently overlooked security indicators, lacked comprehension of their functionality, and found anti-phishing web indicators to be ineffective. Following Dhamija et al.’s research, several similar user studies emerged, featuring participant mock scenarios (downs2006decision; downs2007behavioral; sheng2010falls), where individuals assumed fictional roles.
In a phishing detection user study, Alsharnouby et al. (alsharnouby2015phishing) discovered that participants primarily relied on website appearance, a method proven to be flawed (das2019all). Downs et al. (downs2006decision) identified that people recognized the use of social engineering in phishing attacks, but lacked an understanding of the associated risks. Neupane et al. (neupane2015multi) conducted a multimodal phishing detection study using brain imaging technologies. They presented real and fake website screenshots to users, revealing that users subconsciously process real sites differently from fake sites.
Lastly, research has focused on the role demographics play in phishing detection. Wash (wash2020experts) found that IT experts utilize a three stage strategy to identify phishing attacks. Baki et al. (baki2022sixteen) found that females are at a higher risk of falling for phishing attacks than males. They also found that the elderly outperformed their younger counterparts in detecting fraudulent emails and websites. Similarly, Luga et al. (iuga2016baiting) found that men were more likely to successfully detect phishing attempts than women.
@@
In a smishing study, Rahman et al. (rahman2023users) mentioned that even responding to messages with a request to cease communications can provide a smishing attacker with valuable information. It shows that the phone number is active and the target is willing to engage with suspicious messages. The authors witnessed participants who were willing to interact with phishing messages out of curiosity, even knowing the messages were fraudulent. This can lead to an increased chance of receiving these fraudulent messages and opening users up to future attacks. Smishing research has also focused on aspects such as the evaluation of existing technologies. Timko et al. (timko2023commercial) investigated how effective select bulk messaging services, carriers, and anti-smishing apps were at blocking smishing using a pool of 20 real and fake messages. They found current anti-smishing tools ineffective to protect against modern threats.
The existing literature on smishing detection falls short in comprehensively addressing user perceptions. Loxdal et al. (loxdal2021phishing) conducted a phishing detection user study where participants used the browser app from within a smartphone instead of a computer. They found those who focused on the URL were more likely to correctly determine legitimacy than those who focused on other aspects of the site. As far as we have ascertained, no study has investigated human factors in smishing attacks. Yeboah-Boateng et al. (yeboah2014phishing) conducted a study in which researchers asked users about their perceptions of not just smishing, but also phishing and vishing. The study found that the threat of smishing was between extremely low to moderate amongst their participants.
@@ 3. Background
To measure the effect of different parts of a message on participant decisions, we defined areas of interest (AOI) in each message image and associated user data with clicks in those areas. Alsharnouby et al.’s work (alsharnouby2015phishing) explored grouping user eye tracking data on phishing websites called AOI and measuring how they affect a participant’s ability to successfully identify phishing and legitimate websites. Participants’ performance scores were attributed to these AOI to measure how attention to these areas affected their decision-making.
We utilized research by Rahman et al. (rahman2023users) to identify similar AOI in SMS messages. In this work, differences in message attributes, including the sender, entity, method, call to action, and scenario, were compared to determine their impact on susceptibility to smishing attacks. Here, the sender is the email, phone number, or short code that sent the message. The entity is defined as the perceived organization or brand identified in the message content as the message originator. The method is how the user is asked to respond or interact with the message. The call to action is what the message is asking the user to do. Finally, the scenario refers to the content of the message that motivates the user to take some action.
@@ 4. Methodology
In this section, we present an overview of our study methodology, encompassing ethical considerations, recruitment, demographics, and an explanation of our online survey. A total of 214 smartphone users participated in our study, which comprised of a first round and second round with multiple-choice questions. In the first round task, participants were not informed of the smishing related aspect of our study. The second round involved the smishing detection task, where participants assessed the legitimacy of SMS screenshots.
4.1. Bias, Ethics And Data Protection
To ensure the ethical conduct of our study, we obtained approval from our university’s Institutional Review Board (IRB). At the survey’s outset, participants were informed of potential risks through a consent form. Prior to engaging in the first round task, participants had to explicitly agree to the outlined conditions in the consent form.
To mitigate bias in our study results, we provided participants with an incomplete disclosure regarding the study’s nature. Participants were informed that they would be answering questions related to SMS messages without revealing the study’s security-related focus. In a realistic scenario, information about whether a message is potentially smishing is not disclosed, and such knowledge could potentially sway participants’ choices. Awareness of security or privacy risks might lead participants to exercise more caution or skepticism than they would in everyday situations. Additionally, to measure user response to message content without external influences, we intentionally excluded external context from the SMS, such as whether the message was expected or triggered by a specific activity. Our study’s aim is to gauge reactions to the message content alone, and incomplete disclosure was employed to ensure as accurate results as possible.
In safeguarding participant data against unauthorized access, all personally identifiable information, including survey responses, was stored in encrypted files. Adhering to our university’s IRB guidelines, only members of our research team had access to these files. The consent form assured participants that, while results and findings would be publicized, personally identifiable information (names and email addresses) collected with their consent would remain confidential. Email addresses were solely used for the distribution of opportunity drawing incentive payments. Participation in our study was entirely voluntary, with individuals retaining the right to withdraw consent at any point without repercussion; at withdrawal, all personally identifiable data collected was promptly destroyed. Data from participants who completed the entire study will be securely stored in encrypted files for three years, after which all data will be permanently deleted.
@@ 4.2. Recruitment and Demographics
Our recruitment process involved the distribution of advertising flyers through personal and professional social networks, spanning platforms including Reddit, Twitter, and LinkedIn, in addition to school and personal email groups. The primary goal was to attract participants from the general US population, with the prerequisite that they be over 18 years old and smartphone users. Specifically targeting smartphone users ensured familiarity with SMS messaging. Using Qualtrics as our survey host, we leveraged its features to screen out participants without US IP addresses. As an incentive for study participation, we offered individuals the opportunity to win a \$50 gift card through an optional opportunity drawing. Participants were asked to provide an email address for contact purposes in case they won.
@@ 4.3. Real and Fake Messages Used
The group of SMS screenshots used in the study was selected based on the rate of brand imitation through phishing. Out of the eleven brands used, five were identified as among the most imitated brands in phishing attacks (checkpoint2022dhl). Conversely, we included one brand that was not well-known, chosen from the personal network of a member of our research group. About half of the screenshots originated from the personal networks of our research group members, while the other half originated from SmishTank (smishtankweb; timko2023commercial). Messages were collected from various phones and included different sender types, such as short codes, emails, and phone numbers.
To better represent the current landscape of smishing attacks, we opted to include a screenshot of a romance scam. Romance scams pose a significant threat to users and constitute a multi-million dollar industry in the US (FTC2023). In total, the survey included nine real and seven fake screenshots. Following a precedent set by prior studies (loxdal2021phishing; dhamija2006phishing; alsharnouby2015phishing), we used an uneven number of real and fake examples. Participants were not informed about the total number of real and fake messages in advance.
@@ 4.4. Main Survey
The survey was conducted using Qualtrics, and participants were required to provide consent to participate and commit to offering thoughtful answers. In the first task, participants were presented with an SMS screenshot and asked five questions about the image. This five-question process was repeated 16 times. The complete survey procedure is illustrated in Figure LABEL:fig:mainsurvey in the appendix. Next, we will discuss each of the questions used in our first and second round in the study.
@@ 4.4.1. First Round Questions
Question 1 - Interaction Rate. Initially, participants were asked to rate their likelihood of interacting with the given message on a six-point Likert Scale from ‘Definitely’ to ‘Definitely not.’ A similar user study on phishing emails (VISHWANATH2011576) measured participants’ likelihood to respond to a phishing. Previous research has highlighted that there is some nuance between a user’s likelihood of interacts with messages and their believe that it is fraudulent. For example, some users in phishing studies have noted that, out of curiosity, they willingly interacted with a message even though they believed it was phishing (rahman2023users).
Question 2 - Comfortability Level. Subsequently, participants were asked, ”How comfortable would you feel interacting with this SMS?” on a five-point Likert scale from ‘very comfortable’ to ‘very uncomfortable,’ with an option to select ‘prefer not to say.’ This question was inspired by a previous study on security indicators on websites (Thompson2019), where users were shown images of login pages and asked to rate their comfort with logging in. While there should be some overlap between the answers to this question and the their willingness to interact, we anticipate some differences. Participants who perceive a message as fake may paradoxically feel more comfortable interacting with it, believing they know what to look out for and won’t compromise their information.
Question 3 - AOI Heatmap. We integrated a heat map into our survey, instructing participants to click on message screenshots to indicate the locations affecting their comfort with interacting. This approach, aimed at identifying areas of interest to measure participant focus in security-related tasks, has been well-documented (Thompson2019; 10.1186/s13673-016-0065-2; loxdal2021phishing). Participants were asked to click between three and five times within each message, and the click positions were recorded. These clicks enable us to compare the effects of different parts of the message, referred to as AOI, on participants. Since participants were not made aware of these AOIs during the survey, clicks outside the AOI should not be interpreted as a lack of attention.
Question 4 - Familiarity with Brand. Next, participants were asked a question about their familiarity with the business brand mentioned in the SMS. In a previous study, participants who were familiar with businesses mentioned in phishing emails were more likely to adopt risky behavior when attempting to accurately identify authentic phishing emails (wang2016overconfidence). Here, we employed a similar question to measure its effect on their ability to distinguish between real and fake SMS. Their familiarity with the brand was assessed on a five-point scale ranging from ’extremely familiar’ to ’not familiar at all.’

### v4

# A Quantitative Study of SMS Phishing Detection

@@ Abstract.
With the booming popularity of smartphones, threats related to these devices are increasingly on the rise. Smishing, a combination of SMS (Short Message Service) and phishing has emerged as a treacherous cyber threat used by malicious actors to deceive users, aiming to steal sensitive information, money or install malware on their mobile devices. Despite the increase in smishing attacks in recent years, there are very few studies aimed at understanding the factors that contribute to a user’s ability to differentiate real from fake messages.
To address this gap in knowledge, we have conducted an online survey on smishing detection with 187 participants. In this study, we presented them with 16 SMS screenshots and evaluated how different factors affect their decision making process in smishing detection. Next, we conducted a post-survey to garner information on the participants’ security attitudes, behavior and knowledge. Our results highlighted that attention and security behavioral scores had a significant impact on participants’ accuracy in identifying smishing messages. We found that participants had more difficulty identifying real messages from fake ones, with an accuracy of 67.1% with fake messages and 43.6% with real messages. Our study is crucial in developing proactive strategies to encounter and mitigate smishing attacks. By understanding what factors influence smishing detection, we aim to bolster users’ resilience against such threats and create a safer digital environment for all.
SMS Phishing, Smishing, Usable Security
@@ 1. Introduction
Phishing stands out as one of the most prevalent social engineering attacks in cybersecurity, with Cisco reporting that phishing accounts for 90% of data breaches in the U.S. (cisco, ). These attacks collectively lead to billions of dollars in losses annually (whitehouse, ). Attackers are now extending their efforts into the mobile domain, introducing ransomware, spyware, or adware onto victims’ mobile devices through phishing (bitdefender, ). This malicious software can provide attackers access to sensitive information, including passwords, credit card details, location data, and social security numbers, resulting in significant financial damage (nagunwa2014behind, ).
@@
A survey by Openmarket reveals that 75% of Millennials prefer texting over phone calls, and 83% read SMS messages within 90 seconds of receipt (eztextingstats, ). With the ubiquity of smartphones in daily life, attackers are targeting billions of users (statista, ). RoboKiller’s data indicates a surge in spam texts, with over 225 billion sent in 2022, leading to losses exceeding 22 billion dollars (robocall, ). This represents a 157% increase from the previous year and a staggering 307% rise since 2020. According to the Federal Trade Commission, in 2022, consumers reported losing \$330 million to text message scams, which is more than double the amount reported in 2021 (FTC300million, ).
@@
Numerous classical studies (dhamija2006phishing, ; sheng2010falls, ; alsharnouby2015phishing, ) have sought to understand what makes users vulnerable to email-based phishing attacks. However, research in the realm of smishing remains limited. In a recent study by Rahman et al., (rahman2023users, ), an empirical examination of smishing was carried out. They sent eight fake SMSes to 265 users to gauge the effectiveness of smishing attacks. Their findings revealed that up to 16.92% of the recipients could have fallen victim to these attacks. In a separate study on smishing, Blancaflor et al. (blancaflor2021let, ) discovered that one in 24 targeted users clicked on the phishing URLs within the SMS. These investigations delved into the number of fake SMSes that reached participants, considering that smishing messages could be intercepted by SMS gateways, mobile carriers, or anti-smishing apps on users’ devices (timko2023commercial, ).
@@
To gain a comprehensive understanding of the issue, a more controlled experiment is needed—one that evaluates users’ ability to detect smishing and examines their interactions with both real and fake messages. While previous studies (rahman2023users, ; blancaflor2021let, ) have investigated smishing from the perspective of whether a user actively becomes a victim, we intend to approach the topic from the users’ perception of the messages. We explore what triggers users’ suspicion that a message may be fraudulent, thus causing them to lose their comfort or willingness to interact with it. Additionally, there is nuance between a user actively interacting with a phishing message and whether they believe it is legitimate. End users may perceive a phishing message as legitimate or a legitimate message as fraudulent, while still refraining from interacting with it. By investigating this nuance, we can build a broader picture of smishing and, in turn, help focus security experts and developers on the factors causing end users to make these classification errors.
@@
In this paper, our objective is to bridge the existing gaps by undertaking a comprehensive online study. This will offer a more profound insight into the influence of messaging content and user characteristics on interaction rates, comfort levels, and judgement between real and fake messages. In the main part of the study, we carried out an online user survey focused on smishing detection. To draw correlations with various factors, we incorporated a post-survey that encompasses several standardized metrics.
@@
    We designed and conducted an online survey with 187 participants to gain insights into how different user and message factors affect users’ likelihood to interact, comfort levels with interacting and their judgement of real and fake messages.
@@ 2. Related Works
Research by Dhamija et al. (dhamija2006phishing, ) marked the first investigation into understanding the most effective strategies for deceiving victims in phishing attacks. The study revealed that people consistently overlooked security indicators, lacked comprehension of their functionality, and found anti-phishing web indicators to be ineffective. Following Dhamija et al.’s research, several similar user studies emerged, featuring participant mock scenarios (downs2006decision, ; downs2007behavioral, ; sheng2010falls, ), where individuals assumed fictional roles.
In a phishing detection user study, Alsharnouby et al. (alsharnouby2015phishing, ) discovered that participants primarily relied on website appearance, a method proven to be flawed (das2019all, ). Downs et al. (downs2006decision, ) identified that people recognized the use of social engineering in phishing attacks, but lacked an understanding of the associated risks. Neupane et al. (neupane2015multi, ) conducted a multimodal phishing detection study using brain imaging technologies. They presented real and fake website screenshots to users, revealing that users subconsciously process real sites differently from fake sites.
Lastly, research has focused on the role demographics play in phishing detection. Wash (wash2020experts, ) found that IT experts utilize a three stage strategy to identify phishing attacks. Baki et al. (baki2022sixteen, ) found that the elderly outperformed their younger counterparts in detecting fraudulent emails and websites. Similarly, Luga et al. (iuga2016baiting, ) found that men were more likely to successfully detect phishing attempts than women.
@@
In a smishing study, Rahman et al. (rahman2023users, ) mentioned that even responding to messages with a request to cease communications can provide a smishing attacker with valuable information. It shows that the phone number is active and the target is willing to engage with suspicious messages. The authors witnessed participants who were willing to interact with phishing messages out of curiosity, even knowing the messages were fraudulent. This can lead to an increased chance of receiving these fraudulent messages and opening users up to future attacks. Smishing research has also focused on aspects such as the evaluation of existing technologies. Timko et al. (timko2023commercial, ) investigated how effective select bulk messaging services, carriers, and anti-smishing apps were at blocking smishing using a pool of 20 real and fake messages. They found current anti-smishing tools ineffective to protect against modern threats.
The existing literature on smishing detection falls short in comprehensively addressing user perceptions. Loxdal et al. (loxdal2021phishing, ) conducted a phishing detection user study where participants used the browser app from within a smartphone instead of a computer. They found those who focused on the URL were more likely to correctly determine legitimacy than those who focused on other aspects of the site. Yeboah-Boateng et al. (yeboah2014phishing, ) conducted a study in which researchers asked users about their opinions and perceptions of not just smishing, but also phishing and vishing. The study found that most users had a low level of concern for the threat of smishing, despite 15% of their participants being victims of smishing scams.
@@ 3. Background
To measure the effect of different parts of a message on participant decisions, we defined areas of interest (AOI) in each message image and associated user data with clicks in those areas. Alsharnouby et al.’s work (alsharnouby2015phishing, ) explored grouping user eye tracking data on phishing websites called AOI and measuring how they affect a participant’s ability to successfully identify phishing and legitimate websites. Participants’ performance scores were attributed to these AOI to measure how attention to these areas affected their decision-making.
We utilized research by Rahman et al. (rahman2023users, ) to identify similar AOI in SMS messages. In their work, variations within message attributes, including the sender, entity, method, call to action, and scenario, were compared to determine their impact on susceptibility to smishing attacks. Here, the sender is the email, phone number, or short code that sent the message. The entity is defined as the perceived organization or brand identified in the message content as the message originator. The method is how the user is asked to respond or interact with the message. The call to action is what the message is asking the user to do. Finally, the scenario refers to the content of the message that motivates the user to take some action. Their results showed that some entities and user actions lead to significantly different rates of users falling for smishing attacks. Comparatively, in this current study we look at between the effects of different message attributes on the success rate of smishing attempts.
@@ 4. Methodology
In this section, we present an overview of our study methodology, encompassing ethical considerations, recruitment, demographics, and an explanation of our online survey. A total of 187 smartphone users participated in our study, which comprised of a first round and second round with multiple-choice questions. In the first round task, participants were not informed of the smishing related aspect of our study. The second round involved the smishing detection task, where participants assessed the legitimacy of SMS screenshots.
4.1. Ethical Consideration And Mitigating Biases
To ensure the ethical conduct of our study, we obtained approval from our university’s Institutional Review Board (IRB). At the survey’s outset, participants were informed of potential risks through a consent form. Prior to engaging in the first round task, participants had to explicitly agree to the outlined conditions in the consent form. Participation in our study was entirely voluntary, with individuals retaining the right to withdraw consent at any point without repercussion; at withdrawal, all personally identifiable data collected was promptly destroyed.
To mitigate bias in our study results, we provided participants with an incomplete disclosure regarding the study’s nature. Participants were informed that they would be answering questions related to SMS messages without revealing the study’s security-related focus. Our flyer framed the goal of our research as ”Understand Mobile SMS User Behavior”. This was done to mitigate the response bias, as introducing the topic of smishing may cause the participants to alter their perception of the message or respond more cautiously than in a natural setting.
@@ 4.2. Recruitment and Demographics
Our recruitment process involved the distribution of advertising flyers through personal and professional social networks, spanning platforms including Reddit, Twitter, and LinkedIn,in addition to school email groups. Specifically targeting smartphone users ensured familiarity with SMS messaging. Using Qualtrics as our survey host, we leveraged its features to screen out participants without US IP addresses. As an incentive for participating in the study, we conducted an optional opportunity drawing and gave each of the 12 individuals a \$50 Amazon gift card. Participants were asked to provide an email address for contact purposes in case they won. This process resulted in a total of 187 participants recruited for our study. After recruitment, we performed a post hoc power analysis using GPower (gpower, ) to determine if we had sufficient power to detect medium and large effect sizes. With an alpha of .05 the results showed that we are able to achieve over .8 power for detecting medium and large effect sizes for our within group analysis.
@@ 4.3. Real and Fake Messages Used
The group of SMS screenshots used in the study was selected based on the rate of brand imitation through phishing. Out of the eleven brands used, five were identified as among the most imitated brands in phishing attacks (checkpoint2022dhl, ). Conversely, we included one brand that was not well-known, chosen from the personal network of a member of our research group. These smishing messages originated from SmishTank (smishtankweb, ; timko2023commercial, ), where messages are collected from various phones and included different sender types, such as short codes, emails, and phone numbers. The full list of SMS used in this study can be found in our appendix in Table 6.
To better represent the current landscape of smishing attacks, we opted to include a screenshot of a romance scam. Romance scams pose a significant threat to users and constitute a multi-million dollar industry in the US (FTC2023, ). In total, the survey included nine real and seven fake screenshots. Following a precedent set by prior studies (loxdal2021phishing, ; dhamija2006phishing, ; alsharnouby2015phishing, ), we used an uneven number of real and fake examples. The larger proportion of real messages is more likely to apply to real-world situations, which enhances the ecological validity of our experiment. Participants were not informed about the total number of real and fake messages in advance.
@@ 4.4. Main Survey
The survey was conducted using Qualtrics, and participants were required to provide consent to participate and commit (commitqualtrics, ) to offering thoughtful answers. In the first task, participants were presented with an SMS screenshot and asked five questions about the image. This five-question process was repeated 16 times in randomized order for each participant. The complete survey procedure is illustrated in Figure 2 in the appendix. Next, we will discuss each of the questions used in our first and second round in the study.
@@ 4.4.1. First Round Questions
Question 1 - Interaction Rate. Initially, participants were asked to rate their likelihood of interacting with the given message on a six-point Likert Scale from ‘Definitely’ to ‘Definitely not.’ Interacting with the message was defined as following the URL link, replying to the text or calling the number in the message. A similar user study on phishing emails (VISHWANATH2011576, ) measured participants’ likelihood to respond to a phishing. Previous research has highlighted that there is some nuance between a user’s likelihood of interacts with messages and their belief that it is fraudulent. For example, some users in phishing studies have noted that, out of curiosity, they willingly interacted with a message even though they believed it was phishing (rahman2023users, ).
Question 2 - Comfortability Level. Subsequently, participants were asked, ”How comfortable would you feel interacting with this SMS?” on a five-point Likert scale from ‘very comfortable’ to ‘very uncomfortable,’ with an option to select ‘prefer not to say.’ This question was inspired by a previous study on security indicators on websites (Thompson2019, ), where users were shown images of login pages and asked to rate their comfort with logging in. While there should be some overlap between the answers to this question and the their willingness to interact, we anticipate some differences. Participants who perceive a message as fake may paradoxically feel more comfortable interacting with it, believing they know what to look out for and won’t compromise their information.
Question 3 - AOI Heatmap. We integrated a heat map into our survey, instructing participants to click on message screenshots to indicate the locations affecting their comfort with interacting. This approach, aimed at identifying areas of interest to measure participant focus in security-related tasks, has been well-documented (Thompson2019, ; 10.1186/s13673-016-0065-2, ; loxdal2021phishing, ). Participants were asked to click between three and five times within each message, and the click positions were recorded. These clicks enable us to compare the effects of different parts of the message, referred to as AOI, on participants. Bounding rectangles were drawn around the AOI after the survey concluded and clicks within these rectangles were marked as clicks on the AOI. Since participants were not made aware of these AOIs during the survey, clicks outside the AOI should not be interpreted as a lack of attention.
Question 4 - Familiarity with Brand. Next, participants were asked a question about their familiarity with the business brand mentioned in the SMS. In a previous study, participants who were familiar with businesses mentioned in phishing emails were more likely to adopt risky behavior when attempting to accurately identify authentic phishing emails (wang2016overconfidence, ). Here, we employed a similar question to measure its effect on their ability to distinguish between real and fake SMS. Their familiarity with the brand was assessed on a five-point scale ranging from ’extremely familiar’ to ’not familiar at all.’